OpenStudy (anonymous):

Can you help me: 4.2(y-4)-0.6y=13.2

OpenStudy (anonymous):

1) Distribute the 4.2 through the parentheses 2) Collect (combine) like terms - i.e.: add the coefficient of the terms with the variable, combine the constant terms (no variable) terms seperately. 1): \[4.2y - 16.8 + (-0.6y) = 13.2\] <-- I'd rather add a positive and negative then try to subtract, too easy to make a mistake. 2): \[3.6y = 13.2 + 16.8 -> 3.6y = 30\] \[y = 30/3.6\]